OpenStudy (anonymous):

44xz+53x= find x and z

OpenStudy (anonymous):

the way you wrote it, this problem is impossible to solve .

OpenStudy (anonymous):

tht is the problem on my quiz

OpenStudy (anonymous):

Well then.... is it supposed to be like this, or is it actually equal to something?

OpenStudy (anonymous):

thx for the help man

OpenStudy (anonymous):

How have I helped you ? I didn't.

OpenStudy (anonymous):

44xz+53x=0 x(44z+53)=0\ x=0 44z+53=0 44z=-53 z=-53/44